OpenStudy (anonymous):

Fun Land theme park has a total of 110 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park? A. 80 B. 70 C. 30 D. 20

OpenStudy (lgbasallote):

let x = number of males y = number of females it says the total is 110 so x + y = 110 the number of men is 20 more than twice the number of females so x is 20 more than twice the number of y x = 20 + 2y now substitute this "x" into that other x x + y = 110 20 + 2y + y = 110 20 + 3y = 110 3y = 110 - 20 3y = 90 can you continue now?

OpenStudy (lgbasallote):

note: this is just y

OpenStudy (lgbasallote):

you're looking for x

OpenStudy (anonymous):

y=30

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

so to find x x + 30 = 110

OpenStudy (lgbasallote):

solve for x now

OpenStudy (anonymous):

110-30=x 80=x